Gredvig, Inc. produces and sells a single product. Data concerning that product appear below

Selling price per unit$ 130.00

Variable expense per unit$ 41.60

Fixed expense per month$ 109,616

The break-even in monthly dollar sales is closest to: (Round your intermediate calculations to 2 decimal places.)

$109,616

$161,200

$342,550

$204,455
Answer #1

Contribution margin=Sales-Variable cost  

=(130-41.6)=$88.4

Contribution margin ratio=Contribution margin/Sales

=88.4/130

=0.68

Breakeven=Fixed expenses/Contribution margin ratio

=109,616/0.68

=$161200